OpenStudy (anonymous):

Which of the following equations are identities? More than one.... A. tan^2 x = sec^2 x - 1 B. sin^2 x - cos^2 x = 1 C. sin^2 = 1 - cos^2 x D. cot^2 x + csc^2 x = 1

OpenStudy (anonymous):

you should know sin^2(x) + cos^2(x) = 1 divide both sides by cos^2(x) and you get sin^2(x)/cos^2(x) + 1 = 1/cos^2(x) this becomes tan^2(x) + 1 = sec^2(x) so tan^2(x) = sec^2(x) -1

OpenStudy (anonymous):

sin^2(x) + cos^2(x) = 1 sin^2(x) = 1 - cos^2(x)

OpenStudy (anonymous):

.....?

OpenStudy (anonymous):

...basically a and c are identtiies

OpenStudy (anonymous):

just those? ..for sure?

OpenStudy (anonymous):

yea]